OpenStudy (anonymous):

Write the equation of the line passing through the points (-6,1) and (-4,2).

OpenStudy (nory):

Find the slope of the line. Remember (change in y)/(change in x).

OpenStudy (aravindg):

two point form !

OpenStudy (anonymous):

i really have no idea at all how to find slope or anything graph related my subject is more science and writing lol my teacher actually just gives us the answers and he just gave me this test to do and i honestly dont want his cheat i wan someone to explain it all to me if its no burden

OpenStudy (anonymous):

(y2 - y1) / (x2 - x1).......set 1 (-6,1)....set 2 (-4,2) now sub in your points... (2 - 1)/ (-4 -(-6) 1/2 < -- slope y = mx + b 1 = 1/2(-6) + b 1 = -3 + b 4 = b so your equation is : y = 1/2x + 4

OpenStudy (anonymous):

okay could you explain on how you got all that please ^

OpenStudy (anonymous):

yes... of course First you need to find the slope....this can be done with the formula (y2 - y1) / (x2 - x1).....and you need to label your points as set 1 and set 2 in order to sub these in to the formula. Do you understand this ?

OpenStudy (anonymous):

kind of, so what im getting is you need to first mark those point on the graph?

OpenStudy (anonymous):

no....let me show you how I did this...these are your points (-6,1) this I called set 1....and (-4,2) these I called set 2. Now the formula....m = (y2 - y1)/(x2 - x1).....y2 is the y in set 2 and y1 is the y in set 1. So your first part of the equation is (y2 - y1) which is 2 - 1.....the same goes for x2 and x1.....x2 - x1 is -4 - (-6). Does that help explain it better ?

OpenStudy (anonymous):

yes it does im going to try to apply this to my next problem and if i get stuck ill come back thank you so much <3 :)

OpenStudy (anonymous):

that is just finding the slope....not the total equation...do you need me to continue explaining ?

OpenStudy (anonymous):

oh okay yes please

OpenStudy (anonymous):

By the way, the m in the formulas stand for the slope. Now you need this formula : y = mx + b....m is the slope you found, you can sub either set of points in for x and y, and you just have to find b. Does this make sense so far ?

OpenStudy (anonymous):

okay so after solving y=mx+b i have no to find b right ? okay makes sense

OpenStudy (anonymous):

y = mx + b...we know the m is the slope which we just found changing the equation to : y = 1/2x + b. You following me ? Next we can use either set of points...(-6,1) or (-4,2) either one will work. Lets choose (-6,1)....with -6 being x and 1 being y. now lets sub.. y = 1/2x + b 1 = 1/2(-6) + b 1 = -3 + b 1 + 3 = b 4 = b Now make the equation without the points...y = 1/2x + 4 Does that make sense or do you have any questions ?

OpenStudy (anonymous):

try this....once you put in your given information, it is pretty good about explaining it. http://www.webmath.com/equline1.html

OpenStudy (anonymous):

makes sense thanks, i now know how to do it my self instead of being the pretty girl that doesnt know anything ill go ahead now and look at the site you posted

OpenStudy (anonymous):

you can do this....you sound smart and it is good that you want to learn it instead of just getting answers. If you need any help, just flag me down and i will see what I can do. Good-luck :)

OpenStudy (anonymous):

I wouldn't use this website just to cheat...maybe just check your answers. It is pretty good about explaining